# HiRAG: Retrieval-Augmented Generation with Hierarchical Knowledge

### v2

@@ Appendix D Implementation Details of HiRAG
Appendix E The Clustering Coefficients of HiIndex
We calculate and compare the clustering coefficients of GraphRAG, LightRAG and HiRAG in Figure 5. It evaluates the average connectivity among the neighbors of each entity of our hierarchical KG using the global clustering coefficient
which represents how frequently entities form triangles (closed triples). As shown in Figure 5, HiRAG demonstrates a higher clustering coefficient than other baseline methods, which means that more entities in the hierarchical KG constructed by the HiIndex module tend to cluster together. And the hierarchical KG maintains a proper hierarchy without becoming an overly connected clique (i.e., C=1) or a near-clique. These are the reasons why the HiIndex module can improve the performance of RAG systems.
Appendix F A Simple Case of Hierarchical KG
@@
Appendix G Cross Validation for LLM as a Judge
Although our LLM-based evaluation approach for QFS tasks is a common practice in the performance evaluation by existing graph RAG methods Es et al. [2024] Guo et al. [2024] Edge et al. [2024], we also conducted cross-verification using Qwen-turbo and Claude-3.5-sonnet as the LLM judge to further make our experimental results more convincing. As shown in Table 6, the results consistently demonstrate HiRAG’s superiority over all the other graph RAG methods compared, confirming that our conclusions remain stable even when neutralizing LLM evaluator-specific biases. To reduce the cost, we report the results on the Mix dataset while the results on the other datasets follow a similar pattern.
| Mix | CS | Legal | Agriculture | |
| Recall (%) | Recall (%) | Recall (%) | Recall (%) | |
| Global-level Info | 38.61 | 48.96 | 53.44 | 50.75 |
| Local-level Info | 83.07 | 81.88 | 78.13 | 64.47 |
Appendix H Knowledge Coverage of Bridge-Level Descriptions
To further validate that both local- and global-level knowledge are well connected in the bridge-level descriptions., we counted the average token-level recall ratio of local- and global-level context in the bridge-level context across four QFS datasets. As shown in Table 7, the results demonstrate that our bridge-level retrieval effectively captures a significant portion of both entity-level and community-level information, providing empirical support for the method’s effectiveness. Here, the recall indicates that the percentage of global-level info or local-level info that is captured in the bridge-level context.
Appendix I Prompt Templates used in HiRAG
I.1 Prompt Templates for Entity Extraction
@@
I.2 Prompt Templates for Relation Extraction
@@
I.3 Prompt Templates for Entity Summarization
@@
I.4 Prompt Templates for RAG Evaluation

### v3

@@ Appendix C Implementation Details of HiRAG
Appendix D The Clustering Coefficients of HiIndex
We calculate and compare the clustering coefficients of GraphRAG, LightRAG and HiRAG in Figure 5. HiRAG shows a higher clustering coefficient than other baseline methods, which means that more entities in the hierarchical KG constructed by the HiIndex module tend to cluster together. And this is also the reason why the HiIndex module can improve the performance of RAG systems.
Appendix E A Simple Case of Hierarchical KG
Appendix F Cross Validation for LLM as a Judge
Although our LLM-based evaluation approach for the query-focused summarization task is a common practice in the performance evaluation by existing graph RAG methods Es et al. (2024); Guo et al. (2024); Edge et al. (2024), we also conducted cross-verification using Qwen-turbo and Claude-3.5-sonnet as the LLM judge to further make our experimental results more convincing. As shown in Table 6, the results consistently demonstrate HiRAG’s superiority over all the other graph RAG methods compared, confirming that our conclusions remain stable even when neutralizing LLM evaluator-specific biases. To reduce the cost, we report the results on the Mix dataset while the results on the other datasets follow a similar pattern.
Appendix G Knowledge Coverage of Bridge-Level Descriptions
To further validate that both local- and global-level knowledge are well connected in the bridge-level descriptions, we counted the average token-level recall ratio of local- and global-level context in the bridge-level context across four datasets. As shown in Table 2, the results demonstrate that our bridge-level retrieval effectively captures a significant portion of both entity-level and community-level information, providing empirical support for the method’s effectiveness. Here, the recall indicates that the percentage of global-level info or local-level info that is captured in the bridge-level context.
Appendix H Prompt Templates used in HiRAG
H.1 Prompt Templates for Entity Extraction
@@
H.2 Prompt Templates for Entity Extraction
H.3 Prompt Templates for Relation Extraction
@@
H.4 Prompt Templates for Entity Summarization
@@
H.5 Prompt Templates for RAG Evaluation